OpenStudy (anonymous):

Find the 7th term of the geometric sequence whose common ratio is 3/2 and whose first term is 7. Help appreciated. I'll give medals to.

OpenStudy (raden):

use the formula : an = a1*r^(n-1) here n=7, a1=7, ratio(r) = 3/2